Adrienne deposited some money in an account earning 4% simple interest and $3000 more than
this in an account earning 6% simple interest. After 1 year, she earned a total of $390 in interest.
How much did she deposit in each account?

Answer by stanbon(75887)

Adrienne deposited some money in an account earning 4% simple interest and $3000 more than this in an account earning 6% simple interest.
After 1 year, she earned a total of $390 in interest.
How much did she deposit in each account?
---
Equation:
interest + interest = 390
0.04x + 0.06(x+3000) = 390
----
Multiply thru by 100 to get:
4x + 6x + 6*3000 = 39000
10x = 21000
x = $2100 (amt. invested at 4%)
3000+2100 = $5100 (amt. invested at 6%)
===========================================
Cheers,
Stan H.
===================